The acceleration of a particle at any time t β‰₯ 0 is given by π‘Ž = 𝑑𝑣 𝑑𝑑 = 12 cos 2𝑑 𝑑𝑑𝑖 βˆ’ 8 sin 2𝑑 𝑗 + 16 𝑑 π‘˜ , if the velocity v and displacement r are zero at t=0, find 𝑣 π‘Žπ‘›π‘‘ π‘Ÿ at any time.


1
Expert's answer
2021-05-11T07:29:32-0400

Let's find velocity:

"v(t) = \int {a(t)dt} = \int {12\cos 2t} dt\overrightarrow i - \int {8\sin 2t} dt\overrightarrow j + \int {16tdt} \overrightarrow k = \left( {6\sin 2t + {C_1}} \right)\overrightarrow i + \left( {4\cos 2t + {C_2}} \right)\overrightarrow j + \left( {8{t^2} + {C_3}} \right)\overrightarrow k"

By condition, v(0)=0. Then

"6\sin 0 + {C_1} = 0 \Rightarrow {C_1} = 0"

"4\cos 0 + {C_2} = 0 \Rightarrow 4 + {C_2} = 0 \Rightarrow {C_2} = - 4"

"0 + {C_3} = 0 \Rightarrow {C_3} = 0"

Then

"v(t) = 6\sin 2t\overrightarrow i + \left( {4\cos 2t - 4} \right)\overrightarrow j + 8{t^2}\overrightarrow k"

Let's find displacement:

"r(t) = \int {v(t)dt} = \int {6\sin 2t} dt\overrightarrow i + \int {\left( {4\cos 2t - 4} \right)dt} \overrightarrow j + \int {8{t^2}dt} \overrightarrow k = ( - 3\cos 2t + {C_4})\overrightarrow i + \left( {2\sin 2t - 4t + {C_5}} \right)\overrightarrow j + \left( {\frac{8}{3}{t^3} + {C_6}} \right)\overrightarrow k"

By condition, r(0)=0. Then

"- 3\cos 0 + {C_4} = 0 \Rightarrow - 3 + {C_4} = 0 \Rightarrow {C_4} = 3"

"2\sin 0 - 0 + {C_5} = 0 \Rightarrow {C_5} = 0"

"0 + {C_6} \Rightarrow {C_6} = 0"

Then

"r(t) = ( - 3\cos 2t + 3)\overrightarrow i + \left( {2\sin 2t - 4t} \right)\overrightarrow j + \frac{8}{3}{t^3}\overrightarrow k"

Answer: "v(t) = 6\sin 2t\overrightarrow i + \left( {4\cos 2t - 4} \right)\overrightarrow j + 8{t^2}\overrightarrow k" , "r(t) = ( - 3\cos 2t + 3)\overrightarrow i + \left( {2\sin 2t - 4t} \right)\overrightarrow j + \frac{8}{3}{t^3}\overrightarrow k"